OpenStudy (anonymous):

when we want to find the equation of the locus moving point in the from the given point, what formula should we use? this is coordinate geometry topic.

OpenStudy (anonymous):

@jtvatsim can you help me?

OpenStudy (jtvatsim):

hmm... I'm not too familiar with the equation of the locus.

OpenStudy (jtvatsim):

looks like the equation will be some sort of parabola based on what I'm reading.

OpenStudy (jtvatsim):

I've found something online that seems to make sense. I'll post it below.

OpenStudy (anonymous):

great

OpenStudy (jtvatsim):

This is based off of a quote from math.stackexchange: Let (x,y) be the unknown point. It's distance from the y axis is given by |x|. It's distance from the given point (a,b) is given by \[\sqrt{(x-a)^2 + (y-b)^2}\] Since we want both distances to be equal we must have: \[|x| = \sqrt{(x-a)^2 + (y-b)^2}\] In other words, if we get rid of the square root by squaring: \[x^2 = (x-a)^2 + (y-b)^2\]

OpenStudy (jtvatsim):

I imagine that the above could be simplified given actual numbers. For instance, take (1, 2) as the given point, then the equation is \[x^2 = (x-1)^2 + (y-2)^2\] which becomes \[x^2 = x^2 - 2x + 1 + y^2 - 4y + 4\] or you can write \[x = \frac{1}{2}(y^2 -4y + 5)\]

OpenStudy (anonymous):

so, if the question says: find the equation of the locus of the moving point P (x,y) such that it is equidistant to the points C and D. I must use the formula PC = PD \[ \sqrt{(x-x) + (y-y)} = \sqrt{(x-x) + (y-y)}\]

OpenStudy (anonymous):

?? is it like this?

OpenStudy (jtvatsim):

I think we have to use different notation, let me think for a moment. You are on the right track though. :)

OpenStudy (anonymous):

ok ok.

OpenStudy (jtvatsim):

Actually, I feel like that is a different type of question. I'm basically thinking of it this way: 

OpenStudy (jtvatsim):

But we need the point to be equidistant from C and D so certainly it has to look like this